an anova hypothesis test is equivalent to conducting two sample t-test between every pair of samples. if there are three samples, there are three pairs:1-2,1-3, 2-3. how many pairs are there when there are 4 samples?
Answer #1

here number of pairs are seleting 2 value from n pairs

therefore number of pairs if there are n samples =nC2 =n(n-1)/2

hence number of pairs if there are 3 samples=3*(3-1)/2 =3

number of pairs if there are 4 samples 4*(4-1)/2 =6